# Towards Adapting Federated & Quantum Machine Learning for Network Intrusion Detection: A Survey

### v1

@@ VI-D7 CryptoQFL
In [164], Chu et al. proposed CryptoQFL, which is a secure and efficient QFL framework that extends the baseline QFL by optimizing both computation and communication while maintaining strong security guarantees. In this framework, clients perform hybrid quantum-classical training on their local datasets, encoding the computed gradients into quantum states. Gradients are encrypted using a quantum one-time pad (QOTP), and key updates are performed locally, eliminating the need to transmit QOTP keys to the server, thereby reducing communication overhead and enhancing security. To address the large communication cost associated with high-precision gradient transfer, CryptoQFL employs ternary quantization, representing gradient values with three levels (-1, 0, 1) in a cyclic fashion suitable for quantum parameters. For efficient aggregation of these ternary gradients, a fast multi-bit quantum adder is designed that minimizes the use of costly non-Clifford gates, particularly Toffoli (CCX) gates, while eliminating garbage outputs and enabling serial operations on input qubits. These combined strategies significantly reduce computational latency, communication overhead, and resource consumption, allowing scalable, noise-resilient, and secure federated learning for quantum neural networks without compromising model convergence or accuracy.
@@ VIII Open Challenges in IDS
The increase in digitization and online transactions has led to constant threats to the sensitive and personal data of customers. Also increase in IoT applications, in which sensors collect sensitive data has significantly heightened the risk of data breaches and compromises [24]. There are various challenges for Federated Learning (FL) while implementing IDS [31, 24].
Non-IID Data Distribution: In FL, multiple clients (often referred to as edge nodes or devices) train models on their local data, which can vary significantly across different nodes. These edge nodes are typically distributed across diverse environments and collect data specific to their own context. As a result, the data recorded by these clients often exhibits non-Independent and Identically Distributed (non-IID) characteristics [165]. In other words, the data collected by each client may not follow the same distribution, leading to a heterogeneous data environment.
@@
The non-IID nature of the data presents challenges in terms of model convergence, communication efficiency, and privacy [166]. Since the data across devices differs, the gradients or model updates derived from these clients can vary significantly, causing difficulties in aggregating them effectively. In addition, certain clients might have more abundant data, while others may have limited data, further complicating the model’s ability to learn generalized patterns.
Huang et al. [167] proposed a cluster-based method to deal with non-iid data in IDS. Unlike traditional FedAvg, which aggregates updates from individual nodes, this approach introduces cluster servers to group mutually trusted nodes into clusters. Each cluster operates as a single client in the FL framework. Within these clusters, local updates are aggregated before being passed to the central aggregation server, which then performs global model aggregation. This hierarchical structure significantly reduces the variability caused by non-IID data and enhances the scalability and security of the system. By applying this method, the performance of FL in IoT-based IDS is improved, as demonstrated using the IoT-23 dataset. In [168], Ma et al. surveyed various techniques to deal with non-IID data in federated settings. The authors classified the techniques as data-based, model-based, datasets and data settings, algorithm-based and framework-based and provided future directions based on their survey.
The ability of quantum computing to encode classical data into exponentially high-dimensional Hilbert space can find complex patterns in non-linearly separable data. It can help find shared global patterns despite local heterogeneity. Quantum methods can approximate global gradients across clients with heterogeneous data more robustly [169].
@@
Asynchronous communication is a prominent approach, enabling devices to operate independently without synchronization, thereby mitigating the impact of stragglers caused by variability in computation and communication resources [170]. However, bounded-delay assumptions often used in asynchronous methods must be adapted to the unbounded delays typical in federated settings. Active sampling provides another solution by selectively choosing participating devices based on resource availability or statistical representativeness, ensuring efficient aggregation within time constraints [171]. Yet, extending such methods to handle real-time fluctuations in device performance remains an open challenge. Additionally, fault-tolerant mechanisms are critical, as device failures can bias the federated model if the dropped devices represent specific data distributions [172]. Strategies like ignoring failed devices, though simple, risk introducing bias, emphasizing the need for robust, adaptive techniques to balance fairness and efficiency across heterogeneous devices [173].
In [174], Gurung et al. proposed a model-driven quantum federated learning algorithm (mdQFL), a novel framework to deal with device heterogeneity and non-IID data. Devices are grouped based on similarity between model parameters or performance metrics and a device selection method based on a selection criteria is proposed. The proposed framework shows the power of quantum computing in addressing device heterogeneity. Similar works on client grouping hardware-aware techniques can be found in [175, 176].
@@
  • Inference Attacks: These attacks in FL pose a significant threat to privacy, as adversaries can exploit shared gradients to extract sensitive information about participants’ training data. Gradients, derived from private data, are computed using features and errors from deep learning model layers. Attackers can observe updates and infer details such as class representatives, membership information, or specific properties of the training data. In some cases, they can even reconstruct original training samples without prior knowledge of the dataset [177]. These vulnerabilities arise because deep learning models often encode unintended features, allowing adversaries to infer private details by analyzing differences in consecutive model snapshots or aggregated updates. Such attacks highlight the critical need for robust privacy-preserving mechanisms in FL to protect participants’ data.
@@
  • Poisoning Attacks: Poisoning attacks in FL aim to compromise the integrity of the model by altering its behavior in undesirable ways. These attacks can be random, seeking to reduce model accuracy, or targeted, aiming to misclassify specific inputs [178]. Poisoning can occur at two levels: data and model [179]. Data poisoning involves manipulating the training data, either by subtly altering features without changing labels (clean-label attacks) [180] or by flipping labels to create mislabeled examples (dirty-label attacks) [181]. For example, a malicious participant might flip the labels of all 1s to 7s, causing the model to misclassify 1s consistently. Model poisoning, on the other hand, involves directly tampering with local model updates or embedding hidden backdoors, making the model perform adversarial objectives when triggered by specific input patterns. Targeted model poisoning is particularly effective, as even a single malicious update can introduce backdoors [33]. While data poisoning typically requires less sophistication, model poisoning demands higher technical expertise and resources. Both types of poisoning attacks highlight critical vulnerabilities in FL, especially when adversaries have the capability to manipulate training processes.
@@
  • Sybil Attacks: These attacks in FL involve an attacker creating multiple fake identities to amplify their influence on the global model. This makes FL vulnerable, especially since participants can freely join and leave during training. For example, research shows that just two Sybils can trick the model into misclassifying digits like “1” as “7” [33]. Sybils can use similar or different datasets to craft malicious updates and often coordinate to overpower genuine clients. These attacks highlight a significant challenge in securing FL systems against manipulation.
@@
To mitigate attacks in FL, several defense strategies are employed. Data sanitization filters malicious data before training by leveraging techniques like micro-model voting and global filters (e.g., FedDiv) [182]. However, its reliance on direct data access raises privacy concerns, making it suitable only for trusted environments. Anomaly detection identifies irregularities in client behavior or data patterns through statistical or model-based methods, such as auto-encoders, spectral analysis, and visualization tools like VADAF [25]. It helps exclude malicious clients and anomalous data but faces challenges with unsupervised scenarios and data access restrictions. Differential privacy (DP) ensures that individual data contributions remain indistinguishable, protecting sensitive information from leakage [33]. Similarly, Secure Multi-Party Computation (SMPC) enables collaborative training without revealing data to other parties, offering robust privacy preservation [33].
@@
Threats to FL can be greatly minimized by the use of quantum techniques such as Quantum Key Distribution (QKD) [183] which ensure theoretically unbreakable encryption. Quantum Differential Privacy (QDP) [184] techniques are being explored by researchers and can be future alternatives to the classical counterpart.
@@
Communication Overhead: FL faces several challenges related to communication and computational efficiency [185]. FL can face communication bottlenecks due to a huge number of participating devices, network bandwidth, and a limited amount of computational power as compared to central GPUs. While more devices can improve model training due to diverse data availability, managing many devices simultaneously may lead to communication bottlenecks and increased computational costs [37]. FL reduces traditional ML costs, but reliable bandwidth is important. Unstable networks or mismatched upload/download speeds can delay model uploads, disrupting training. Edge devices, unlike powerful GPUs, have limitations in processing power, storage, and bandwidth. Training complex models on edge devices is significantly slower, even with advanced connectivity like 5G [42].
@@
In FL, local updating, client selection, decentralized training and peer-to-peer learning, model update and compression schemes reduction are critical techniques to improve communication efficiency [185]. Briggs et al., in [186], the authors combine FL with Hierarchical Clustering (HC) to reduce communication rounds by grouping clients based on the similarity of their local updates to the global model. This approach, tested using the Manhattan distance metric, reduces the number of communication rounds.
@@
Although classical data compression and asynchronous methods have mitigated some of the communication issues, emerging quantum advancements can bring a radical shift. Techniques such as quantum teleportation [187], and entanglement-assisted methods [188] may enable ultra-efficient and secure transmission of model updates between clients and servers. Furthermore, quantum data compression [189] and quantum gradient updates [190, 191] are being explored that could reduce the communication footprint in QFL.
@@ IX-A Regulatory Compliance
Data Privacy Regulations: Laws such as the General Data Protection Regulation (GDPR) [192] in Europe, the California Consumer Privacy Act (CCPA) [193], and national cybersecurity frameworks impose strict requirements on collection, storage, and sharing of personal and sensitive data. FL partially mitigates privacy risks by keeping raw data on local devices, but mechanisms like differential privacy and secure aggregation are necessary to comply with regulatory standards.
Industry-Specific Standards: Critical infrastructure operators, financial institutions, and healthcare organizations may need to adhere to sector-specific regulations (e.g., NIST Cybersecurity Framework, ISO/IEC 27001) when integrating FL into intrusion detection pipelines [194, 195].
@@ IX-B Ethical Considerations
Transparency and Explainability: Security teams must understand why a model flags specific traffic as malicious. Explainable AI (XAI) [196] methods are critical to maintain trust in automated detection systems.
@@ X Industrial Adoption Roadmap
    Standardized Evaluation & Continuous Improvement: Finally, the fifth phase highlights the use of standardized benchmarks such as UNSW-NB15 [197] and Bot-IoT [198], defines cross-industry IDS benchmarks, and stresses continuous model re-training to handle evolving cyber threats while enabling real-time anomaly detection.
@@ XI Future Directions
To fully realize the potential of FL, several key challenges need attention. Enhancing privacy-preserving techniques is of utmost importance. Approaches like advanced differential privacy mechanisms, zero-knowledge proofs, and blockchain integration [199] can protect sensitive data while maintaining system integrity. For instance, blockchain can ensure the immutability and reliability of model updates [200, 201], while quantum-resistant cryptography [183] can prepare FL for future challenges posed by quantum computing [202].
Optimizing computational and communication efficiency is crucial, especially for large-scale IDS deployments. Novel optimization algorithms that accelerate convergence, reduce the computational burden, and minimize communication overhead are vital. Techniques like selective data sharing, model compression [203], and secure aggregation can address these challenges without compromising security. Current quantum hardware is limited by noise, qubit count, and coherence times. Developing QFL algorithms and error-mitigating techniques for NISQ devices can be a great innovation in the real world IDS deployment. Designing effective quantum feature encoding techniques particularly tailored to capture complex network traffic patterns can enhance IDS.
While QFL provides privacy benefits, quantum-specific threats and vulnerabilities (such as quantum adversarial attacks) [204] require new quantum-safe privacy preservation methods to be developed.

### v2

@@ VI-D7 CryptoQFL
In [chu2023cryptoqfl], Chu et al. proposed CryptoQFL, which is a secure and efficient QFL framework that extends the baseline QFL by optimizing both computation and communication while maintaining strong security guarantees. In this framework, clients perform hybrid quantum-classical training on their local datasets, encoding the computed gradients into quantum states. Gradients are encrypted using a quantum one-time pad (QOTP), and key updates are performed locally, eliminating the need to transmit QOTP keys to the server, thereby reducing communication overhead and enhancing security. To address the large communication cost associated with high-precision gradient transfer, CryptoQFL employs ternary quantization, representing gradient values with three levels (-1, 0, 1) in a cyclic fashion suitable for quantum parameters. For efficient aggregation of these ternary gradients, a fast multi-bit quantum adder is designed that minimizes the use of costly non-Clifford gates, particularly Toffoli (CCX) gates, while eliminating garbage outputs and enabling serial operations on input qubits. These combined strategies significantly reduce computational latency, communication overhead, and resource consumption, allowing scalable, noise-resilient, and secure federated learning for quantum neural networks without compromising model convergence or accuracy.
@@ VIII Open Challenges in IDS
The increase in digitization and online transactions has led to constant threats to the sensitive and personal data of customers. Also increase in IoT applications, in which sensors collect sensitive data has significantly heightened the risk of data breaches and compromises [alazab2021federated]. There are various challenges for Federated Learning (FL) while implementing IDS [agrawal2022federated, alazab2021federated].
Non-IID Data Distribution: In FL, multiple clients (often referred to as edge nodes or devices) train models on their local data, which can vary significantly across different nodes. These edge nodes are typically distributed across diverse environments and collect data specific to their own context. As a result, the data recorded by these clients often exhibits non-Independent and Identically Distributed (non-IID) characteristics [lu2024federated]. In other words, the data collected by each client may not follow the same distribution, leading to a heterogeneous data environment.
@@
The non-IID nature of the data presents challenges in terms of model convergence, communication efficiency, and privacy [li2022federated]. Since the data across devices differs, the gradients or model updates derived from these clients can vary significantly, causing difficulties in aggregating them effectively. In addition, certain clients might have more abundant data, while others may have limited data, further complicating the model’s ability to learn generalized patterns.
Huang et al. [huang2022federated] proposed a cluster-based method to deal with non-iid data in IDS. Unlike traditional FedAvg, which aggregates updates from individual nodes, this approach introduces cluster servers to group mutually trusted nodes into clusters. Each cluster operates as a single client in the FL framework. Within these clusters, local updates are aggregated before being passed to the central aggregation server, which then performs global model aggregation. This hierarchical structure significantly reduces the variability caused by non-IID data and enhances the scalability and security of the system. By applying this method, the performance of FL in IoT-based IDS is improved, as demonstrated using the IoT-23 dataset. In [ma2022state], Ma et al. surveyed various techniques to deal with non-IID data in federated settings. The authors classified the techniques as data-based, model-based, datasets and data settings, algorithm-based and framework-based and provided future directions based on their survey.
The ability of quantum computing to encode classical data into exponentially high-dimensional Hilbert space can find complex patterns in non-linearly separable data. It can help find shared global patterns despite local heterogeneity. Quantum methods can approximate global gradients across clients with heterogeneous data more robustly [qi2023optimizing].
@@
Asynchronous communication is a prominent approach, enabling devices to operate independently without synchronization, thereby mitigating the impact of stragglers caused by variability in computation and communication resources [recht2011hogwild]. However, bounded-delay assumptions often used in asynchronous methods must be adapted to the unbounded delays typical in federated settings. Active sampling provides another solution by selectively choosing participating devices based on resource availability or statistical representativeness, ensuring efficient aggregation within time constraints [nishio2019client]. Yet, extending such methods to handle real-time fluctuations in device performance remains an open challenge. Additionally, fault-tolerant mechanisms are critical, as device failures can bias the federated model if the dropped devices represent specific data distributions [yin2018byzantine]. Strategies like ignoring failed devices, though simple, risk introducing bias, emphasizing the need for robust, adaptive techniques to balance fairness and efficiency across heterogeneous devices [bonawitz2019towards].
In [gurung2025communication], Gurung et al. proposed a model-driven quantum federated learning algorithm (mdQFL), a novel framework to deal with device heterogeneity and non-IID data. Devices are grouped based on similarity between model parameters or performance metrics and a device selection method based on a selection criteria is proposed. The proposed framework shows the power of quantum computing in addressing device heterogeneity. Similar works on client grouping hardware-aware techniques can be found in [10623010, 10758814].
@@
  • Inference Attacks: These attacks in FL pose a significant threat to privacy, as adversaries can exploit shared gradients to extract sensitive information about participants’ training data. Gradients, derived from private data, are computed using features and errors from deep learning model layers. Attackers can observe updates and infer details such as class representatives, membership information, or specific properties of the training data. In some cases, they can even reconstruct original training samples without prior knowledge of the dataset [zhu2019deep]. These vulnerabilities arise because deep learning models often encode unintended features, allowing adversaries to infer private details by analyzing differences in consecutive model snapshots or aggregated updates. Such attacks highlight the critical need for robust privacy-preserving mechanisms in FL to protect participants’ data.
@@
  • Poisoning Attacks: Poisoning attacks in FL aim to compromise the integrity of the model by altering its behavior in undesirable ways. These attacks can be random, seeking to reduce model accuracy, or targeted, aiming to misclassify specific inputs [huang2011adversarial]. Poisoning can occur at two levels: data and model [lyu2020threats]. Data poisoning involves manipulating the training data, either by subtly altering features without changing labels (clean-label attacks) [shafahi2018poison] or by flipping labels to create mislabeled examples (dirty-label attacks) [gu2017badnets]. For example, a malicious participant might flip the labels of all 1s to 7s, causing the model to misclassify 1s consistently. Model poisoning, on the other hand, involves directly tampering with local model updates or embedding hidden backdoors, making the model perform adversarial objectives when triggered by specific input patterns. Targeted model poisoning is particularly effective, as even a single malicious update can introduce backdoors [li2024threats]. While data poisoning typically requires less sophistication, model poisoning demands higher technical expertise and resources. Both types of poisoning attacks highlight critical vulnerabilities in FL, especially when adversaries have the capability to manipulate training processes.
@@
  • Sybil Attacks: These attacks in FL involve an attacker creating multiple fake identities to amplify their influence on the global model. This makes FL vulnerable, especially since participants can freely join and leave during training. For example, research shows that just two Sybils can trick the model into misclassifying digits like “1” as “7” [li2024threats]. Sybils can use similar or different datasets to craft malicious updates and often coordinate to overpower genuine clients. These attacks highlight a significant challenge in securing FL systems against manipulation.
@@
To mitigate attacks in FL, several defense strategies are employed. Data sanitization filters malicious data before training by leveraging techniques like micro-model voting and global filters (e.g., FedDiv) [cretu2008casting]. However, its reliance on direct data access raises privacy concerns, making it suitable only for trusted environments. Anomaly detection identifies irregularities in client behavior or data patterns through statistical or model-based methods, such as auto-encoders, spectral analysis, and visualization tools like VADAF [chen2018autoencoder]. It helps exclude malicious clients and anomalous data but faces challenges with unsupervised scenarios and data access restrictions. Differential privacy (DP) ensures that individual data contributions remain indistinguishable, protecting sensitive information from leakage [li2024threats]. Similarly, Secure Multi-Party Computation (SMPC) enables collaborative training without revealing data to other parties, offering robust privacy preservation [li2024threats].
@@
Threats to FL can be greatly minimized by the use of quantum techniques such as Quantum Key Distribution (QKD) [renner2008security] which ensure theoretically unbreakable encryption. Quantum Differential Privacy (QDP) [hirche2023quantum] techniques are being explored by researchers and can be future alternatives to the classical counterpart.
@@
Communication Overhead: FL faces several challenges related to communication and computational efficiency [shahid2021communication]. FL can face communication bottlenecks due to a huge number of participating devices, network bandwidth, and a limited amount of computational power as compared to central GPUs. While more devices can improve model training due to diverse data availability, managing many devices simultaneously may lead to communication bottlenecks and increased computational costs [mothukuri2021survey]. FL reduces traditional ML costs, but reliable bandwidth is important. Unstable networks or mismatched upload/download speeds can delay model uploads, disrupting training. Edge devices, unlike powerful GPUs, have limitations in processing power, storage, and bandwidth. Training complex models on edge devices is significantly slower, even with advanced connectivity like 5G [khraisat2024survey].
@@
In FL, local updating, client selection, decentralized training and peer-to-peer learning, model update and compression schemes reduction are critical techniques to improve communication efficiency [shahid2021communication]. Briggs et al., in [briggs2020federated], the authors combine FL with Hierarchical Clustering (HC) to reduce communication rounds by grouping clients based on the similarity of their local updates to the global model. This approach, tested using the Manhattan distance metric, reduces the number of communication rounds.
@@
Although classical data compression and asynchronous methods have mitigated some of the communication issues, emerging quantum advancements can bring a radical shift. Techniques such as quantum teleportation [pirandola2015advances], and entanglement-assisted methods [hsieh2010entanglement] may enable ultra-efficient and secure transmission of model updates between clients and servers. Furthermore, quantum data compression [rozema2014quantum] and quantum gradient updates [kerenidis2020quantum, gilyen2019optimizing] are being explored that could reduce the communication footprint in QFL.
@@ IX-A Regulatory Compliance
Data Privacy Regulations: Laws such as the General Data Protection Regulation (GDPR) [GDPR] in Europe, the California Consumer Privacy Act (CCPA) [CCPA], and national cybersecurity frameworks impose strict requirements on collection, storage, and sharing of personal and sensitive data. FL partially mitigates privacy risks by keeping raw data on local devices, but mechanisms like differential privacy and secure aggregation are necessary to comply with regulatory standards.
Industry-Specific Standards: Critical infrastructure operators, financial institutions, and healthcare organizations may need to adhere to sector-specific regulations (e.g., NIST Cybersecurity Framework, ISO/IEC 27001) when integrating FL into intrusion detection pipelines [NIST, ISO].
@@ IX-B Ethical Considerations
Transparency and Explainability: Security teams must understand why a model flags specific traffic as malicious. Explainable AI (XAI) [nwakanma2023explainable] methods are critical to maintain trust in automated detection systems.
@@ X Industrial Adoption Roadmap
    Standardized Evaluation & Continuous Improvement: Finally, the fifth phase highlights the use of standardized benchmarks such as UNSW-NB15 [7348942] and Bot-IoT [koroniotis2019towards], defines cross-industry IDS benchmarks, and stresses continuous model re-training to handle evolving cyber threats while enabling real-time anomaly detection.
@@ XI Future Directions
To fully realize the potential of FL, several key challenges need attention. Enhancing privacy-preserving techniques is of utmost importance. Approaches like advanced differential privacy mechanisms, zero-knowledge proofs, and blockchain integration [pokhrel2020federated] can protect sensitive data while maintaining system integrity. For instance, blockchain can ensure the immutability and reliability of model updates [8843900, xu2022fair], while quantum-resistant cryptography [renner2008security] can prepare FL for future challenges posed by quantum computing [li2024enhancing].
Optimizing computational and communication efficiency is crucial, especially for large-scale IDS deployments. Novel optimization algorithms that accelerate convergence, reduce the computational burden, and minimize communication overhead are vital. Techniques like selective data sharing, model compression [shah2021model], and secure aggregation can address these challenges without compromising security. Current quantum hardware is limited by noise, qubit count, and coherence times. Developing QFL algorithms and error-mitigating techniques for NISQ devices can be a great innovation in the real world IDS deployment. Designing effective quantum feature encoding techniques particularly tailored to capture complex network traffic patterns can enhance IDS.
While QFL provides privacy benefits, quantum-specific threats and vulnerabilities (such as quantum adversarial attacks) [liao2021robust] require new quantum-safe privacy preservation methods to be developed.